OpenStudy (anonymous):

How many solutions does the equation have? 3(4x - 2) = 2(6x - 3)

OpenStudy (lynfran):

solve it to know

OpenStudy (lynfran):

first distribute the parentheses

OpenStudy (lynfran):

so 3*4x-3*2=2*6x-2*3

OpenStudy (anonymous):

so 12x - 2 = 12x - 6?

OpenStudy (lynfran):

12x-6=12x-6

OpenStudy (lynfran):

so we get 0=0

OpenStudy (lynfran):

12x-12x=-6+6 0=0

OpenStudy (lynfran):

we have an infinite many solution

OpenStudy (anonymous):

thank you!

OpenStudy (lynfran):

welcome